OpenStudy (anonymous):

can someone help me to prove that the dimensions of the electric field are: MLT^(-2)Q^(-1)

OpenStudy (anonymous):

Electric field has dimensions of (Force/Unit Charge). Force is Newton = mass*acceleration = kg.m/s^2 = MLT^(-2) Charge is Coulomb = Q Hence Electric field has dimensions MLT^(-2)/Q = MLT^(-2)Q^(-1)

OpenStudy (anonymous):

e = f/q f= m*a = MLT^-2 THUS E= MLT-^2 Q-^1